Question #272778

Find the probability that of the next 200 children born, (a) less than 40% will be boys, (b) between 43% and 57% will be girls, (c) more than 54% will be boys. Assume equal probabilities for births of boys and girls

Expert's answer

If XX is a binomial random variable with mean μ=npμ = np and variance σ2=np(1−p),\sigma^2=np(1-p), then the limiting form of the distribution of

Z=X−npnp(1−p),Z=\dfrac{X-np}{\sqrt{np(1-p)}},

as n→∞,n\to \infin, is the standard normal distribution n(z;0,1).n(z; 0, 1).

In practice, the approximation is adequate provided that both np≥10np\geq10 and np(1−p)≥10.np(1-p)\geq 10.

a) n=200,p=0.5,μ=np=200(0.5)=100,n=200, p=0.5, \mu=np=200(0.5)=100,

σ2=np(1−p)=200(0.5)(1−0.5)=50\sigma^2=np(1-p)=200(0.5)(1-0.5)=50


P(X<80)≈P(X<80−0.5)P(X<80)\approx P(X<80-0.5)

≈P(Z<−2.8991378)\approx P(Z<-2.8991378)

≈0.001871\approx0.001871

b) n=200,p=0.5,μ=np=200(0.5)=100,n=200, p=0.5, \mu=np=200(0.5)=100,

σ2=np(1−p)=200(0.5)(1−0.5)=50\sigma^2=np(1-p)=200(0.5)(1-0.5)=50


P(86<X<114)=P(X<114)−P(X≤86)P(86<X<114)=P(X<114)-P(X\leq 86)

≈P(X<114−0.5)−P(X<86+0.5)\approx P(X<114-0.5)-P(X<86+0.5)

=P(Z<113.5−10050)−P(Z<86.5−10050)=P(Z<\dfrac{113.5-100}{\sqrt{50}})-P(Z<\dfrac{86.5-100}{\sqrt{50}})

≈P(Z<1.90918831)−P(Z<−1.90918831)\approx P(Z<1.90918831)-P(Z<-1.90918831)

≈0.9718811−0.0281189≈0.943762\approx0.9718811-0.0281189\approx0.943762

c) n=200,p=0.5,μ=np=200(0.5)=100,n=200, p=0.5, \mu=np=200(0.5)=100,

σ2=np(1−p)=200(0.5)(1−0.5)=50\sigma^2=np(1-p)=200(0.5)(1-0.5)=50


P(X>108)=1−P(X≤108)P(X>108)=1-P(X\leq 108)

≈1−P(X<108+0.5)\approx 1-P(X<108+0.5)

=1−P(Z<108.5−10050)=1-P(Z<\dfrac{108.5-100}{\sqrt{50}})

≈1−P(Z<1.20208153)\approx1-P(Z<1.20208153)

≈0.114666\approx 0.114666


LATEST TUTORIALS
APPROVED BY CLIENTS